(81k^4m^-8)^1/4
use the laws of exponets to simplify. do not use negitive exponets in your answer.

Answer by fractalier(2101)

Okay, using the laws of exponents...we multiply when raising a power to another power and get
(81k^4m^-8)^1/4 =
81^(1/4) * k * m^(-2) =
3k / m^2